OpenStudy (anonymous):

The population of a certain town increased by 2% from 35,000. What was the population after the increase?

OpenStudy (anonymous):

So, if it increased 2%, it now has a total of 100% that it had before, plus a 2% from the growth for a total of 102%. 102% of 35000 can then be written in the following manner. \[{102 \over 100} \times 35000 = 35700\] Or you could calculate what the 2% are and only then add it to what you began with. \[{2 \over 100} \times 35000 = 700\] 35000 + 700 = 35700.